OpenStudy (anonymous):

convert 17lbs into ounces..................how many ounces will it be

OpenStudy (anonymous):

how many ounces in a pound. you are already on the internet, you could look it up?

OpenStudy (angela210793):

1pound=16ounce 17 pounds=x ounce x=17*16=272 ounce

OpenStudy (anonymous):

1 lb = 16 oz. 17 * 16 = 272 oz